OpenStudy (anonymous):

Sorry I'm still having trouble on these problems What is the value of b^2 - 4ac for the following equation? 2x^2 - 2x - 1 = 0 a) -4 b) 0 c) 12 I get that A= 2 B= -2 C= -1

OpenStudy (anonymous):

it's [b ^{2}-4ac\] But I don't remember why or how. To solve it it's easy, think of your equation in this way: a is paired with x because x is earlier in the alphabet, so is a b is paired with y because y is after x in the alphabet, as is b to a and c is always your number on it's own.

OpenStudy (anonymous):

by basic quadratic equation solving formula For ax2 + bx + c = 0, the value of x is given by: x1=(-b+(b^2-4ac))/2a X2=(-b-(b^2-4ac))/2a x1=0.95 x2=0.786

OpenStudy (anonymous):

So you need to find a,b, and c, and use the above formula the a, b, and c come from a x^2 + bx + c so match your quadratic 2x^2 + 7x + 3 = 0 to find a,b, and c.

OpenStudy (anonymous):

it is c

OpenStudy (anonymous):

ty